Given lê – X-3X,is an estimator of the population mean μ, where X 1and X 2 are from the same distribution with mean μ and standard deviation σ.

1. Find the following:

a.  E ( Θ ^ )

b.  B i a s ( Θ ^ )

c. Is Θ ^  a biased estimator of μ? Please briefly explain your answer.

d.  V ( Θ ^ )

e. MSE ( Θ ^ )
Answer #1

Given that and E(X) = E(X2) = f VC x1) = V(x2) = 02 (0) E (6) - ( xi-342) { EC) – 3 E64239 - + { M-343 -- 2 - - (6) : Bias CôCO (Ⓡ) MSECô) (Biasco)] + v() = 62m² + 5 = 22 - Hep + 50%